OpenStudy (anonymous):

A boy drops a 0.10 kg stone down a 150 m well and listens for the echo. The air temperature is 20°C. How long after the stone is dropped will the boy hear the echo? The speed of sound through air at 20°C is v = 343 m/s.

OpenStudy (shamim):

Formula needed h=ut+0.5gt^2 150=0*t+0.5*9.8*t^2 t=?

OpenStudy (shamim):

The above time is needed to touch the water of well by the stone.

OpenStudy (shamim):

Can u tell me the time needed to travel the sound frm water to the top of well?

OpenStudy (shamim):

t=150/343=?

OpenStudy (anonymous):

Im sorry I really do not understand

OpenStudy (shamim):

Ok. I wanna try it again!!!!

OpenStudy (shamim):

Do u know the formula S=ut+(1/2)at^2

OpenStudy (shamim):

For free fall body u hv to write a=g

OpenStudy (shamim):

So the equation for a free body h=ut+(1/2)gt^2 Right???

OpenStudy (shamim):

Given h=150m u=0 g=9.8m/s^2 t=?

OpenStudy (anonymous):

I'm not familiar with that formula

OpenStudy (shamim):

Ok

OpenStudy (shamim):

So u r not familiar with this formula S=ut+(1/2)at^2 Right?

OpenStudy (anonymous):

That is correct. Im having issues with my internet connection so Im going to go. I appreciate the help though

OpenStudy (shamim):

Medal????!!!

OpenStudy (anonymous):

Hm.. since you're not too familiar with formulas, we can try with concepts, sorry for the wall of text, hope it helps. The stone falls to the bottom of the well, this will take a time t1 Sound is generated and travels to the top, which takes time t2 The equation needed to calculate t1 is the one from linear motion with constant acceleration, which is the one Shamim provided: (change in position) = (initial speed)*(time) + (1/2)*(acceleration)*(time)^2 Since the acceleration is the accel of gravity, which is equals to 9,81m/s^2, and the initial speed is 0, because the rock was dropped, not thrown; and the change in position is equal to the depth of the well, 150m, you replace those values and solve for time. Now, to calculate t2, you use the equation from linear motion at constant speed: (change in position)=(speed)*(time) You can now solve for time replacing the change in position as 150m, and the speed, which is the speed of sound = 343m/s Finally, you add up t1 + t2 and done :D